# Frame-Voyager: Learning to Query Frames for Video Large Language Models

### v2

@@ 1 Introduction
Recent studies (Liu et al., 2023; 2024a; Li et al., 2024c; Lin et al., 2024b) explore integrating Large Language Models (LLMs, Stiennon et al. (2020); Gao et al. (2023); OpenAI (2023); Touvron et al. (2023); Jiang et al. (2023); Yang et al. (2024)) with visual foundation models (e.g., Vision Transformer (ViT, Dosovitskiy et al. (2021)) and cross-modal projectors (Li et al., 2023a; Lin et al., 2024a; Liu et al., 2023). In this paper, we focus on Video-LLMs. Existing Video-LLMs (Zhang et al., 2023; Cheng et al., 2024; Li et al., 2024b) usually treat the video as a sequence of image frames. The key challenge is that the entire video can not be fed into the model due to LLMs’ token length limitation (Xue et al., 2024). Meanwhile, arbitrarily increasing the token length of the model (Miao et al., 2023; Wan et al., 2024; Xiong et al., 2024; Zhang et al., 2024a) may lead to the “lost-in-the-middle” issue (Liu et al., 2024c) and introduce significant computational complexity.
To mitigate this issue, some efforts propose to select only a subset of frames as input, e.g., through uniform sampling (Liu et al., 2023; Lin et al., 2024b; Cheng et al., 2024) or text-frame matching (Liang et al., 2024; Wang et al., 2024; Yu et al., 2024). The uniform sampling strategy evenly samples frames in videos, while text-frame matching typically retrieves a set of relevant frames by calculating semantic similarities, e.g., using CLIP (Radford et al., 2021), between the query and each frame. However, the uniform sampling fails to account for the information density variations in the videos. For instance, in the video question answering task, answering different questions may rely on distinct video segments or frames (Fu et al., 2024a). Meanwhile, text-frame matching is inadequate for complex video understanding tasks that require multi-frame or temporal reasoning, such as tracking the progression of an action or understanding cause-and-effect relationships over time. For instance, in the video summarization task, simply matching frames might overlook the subtle transitions that connect scenes, while in temporal reasoning tasks—such as answering why does the woman need to drink water at the beginning of the video?—it is crucial to concentrate on the beginning part of the video. These matching-based methods fail to account for these frame-to-frame interactions and the relative positional information essential for a comprehensive understanding of the video. To solve these problems, we introduce an innovative approach named Frame-Voyager that learns to query the subset of frames in a combinational manner, rather than retrieving individual frames separately. This capability is essential for understanding dynamic scenes and the global context of events.
@@ 2 Related Work
However, in existing Video-LLM models, a single frame is typically represented by $64$-$256$ visual tokens (Liu et al., 2023; Lin et al., 2024b; Cheng et al., 2024). Due to the input limitations of LLMs, the number of frames that can be processed by Video-LLM models is often constrained. Some studies apply techniques for handling long LLM inputs to support more frames (Miao et al., 2023; Wan et al., 2024; Xiong et al., 2024; Xue et al., 2024; Song et al., 2024), but this approach significantly increases computational complexity and can lead to issues such as the “lost-in-the-middle” effect and hallucinations (Liu et al., 2024c). Other studies, while keeping the frame input limit unchanged, use alternative sampling strategies instead of the default uniform sampling to obtain higher-quality frames as inputs. Some initial attempts focus on identifying transition frames (Lu & Grauman, 2013; Rochan et al., 2018; Rochan & Wang, 2019) or using frame clustering (Liang et al., 2024) to find central frames, but these methods often overlook the information from the query. Subsequent research view this problem as an individual text-frame matching task (Wang et al., 2024; Yu et al., 2024), attempting to select frames that are semantically closest to the query. However, this method is still sub-optimal as it ignores frame-to-frame relationships, failing in complex video understanding tasks that require multi-frame or temporal information. Therefore, to the best of our knowledge, our method is the first to consider the combination of frames as a whole, aiming to find the optimal combination that can best answer the query under the constraint of frame length limitations.
@@ 3.1 Data Collection
Our pipeline is based on a simple intuition: if one frame combination is better than another, it will produce a lower language modeling loss when used as input to any trained Video-LLM. For each (video, query) pair, we evaluate all possible combinations of $T$ frames selected from a total of $M$ video frames, resulting in $\mathcal{C}(M,T)$ combinations, where $\mathcal{C}(M,T)$ is the binomial coefficient representing the number of ways to choose $T$ items from $M$. Each combination, along with the query, is then input into a trained reference Video-LLM to calculate the combination loss, i.e., language modeling loss against the ground-truth answer. We collect the loss values for each combination and rank them from best to worst by sorting the losses in ascending order. It is worth noting that as $M$ increases, the potential number of combinations $\mathcal{C}(M,T)$ grows exponentially, making exhaustive traversal computationally infeasible. For example, when $M=64$ and $T=8$, the total number of combinations is approximately $\mathcal{C}(64,8)\approx 4\times 10^{9}$. Considering that the majority of the training data are with short videos, we use smaller combinations during training, such as $\mathcal{C}(16,2)$ or $\mathcal{C}(32,4)$. We observe from experiments that models trained with smaller combinations exhibit generalization capabilities when larger values of $M$ and $T$ are used during inference for longer video or complex reasoning. The specific choices of $M$ and $T$ employed in our experiments are detailed in Section 4.1.
@@ 3.2 Model Training and Inference
4 Experiments
4.1 Experiment Settings
Backbone Models. We use two variants of the state-of-the-art Video-LLM called VILA (Lin et al., 2024b): VILA-8B and VILA-40B. VILA-8B employs SigLIP (Zhai et al., 2023) as its visual encoder and Llama3-8B (Dubey et al., 2024) as its LLM backbone, while VILA-40B utilizes InternViT-6B (Chen et al., 2024) for visual encoding and Yi-34B (Young et al., 2024) as its LLM backbone.
@@ 4.1 Experiment Settings
+Frame-Voyager 8B 50.5 / 53.6 60.3 / 65.0 47.3 / 50.3 43.9 / 45.3 49.8 51.1 51.6
@@
+Frame-Voyager 34B 60.0 / 63.8 70.3 / 73.1 58.3 / 62.7 51.2 / 55.7 61.1 55.2 69.4
Benchmarks. We evaluate Frame-Voyager on four widely-adopted video benchmarks: Video-MME (Fu et al., 2024a), MLVU (Zhou et al., 2024), NextQA (Xiao et al., 2021) and ActivityNet-QA (Yu et al., 2019). The former two evaluation datasets are tailored for long video assessments, while the latter two focus on short videos. For the former (the latter), we uniformly downsample the video to $128$ ($32$) candidate frames and each time select $8$ ($2$) frames to compose a frame combination. The LMMs-Eval Library (Li et al., 2024a) is used for evaluation, and accuracy is reported across all benchmarks. Note that we report the accuracy scores of Video-MME under both without (no sub.) and with (sub.) subtitles settings.
Implementation Details. We reuses one bottom layer from the LLM for Frame-Voyager in Section 3.2. During the Frame-Voyager dataset construction, VILA-8B is consistently adopted as the reference Video-LLM for generating loss due to resource limitations. During training, we set $K\!=\!4$ for sampling combination ranking data. VILA-8B is trained using DeepSpeed (Aminabadi et al., 2022) ZeRO2 with $8$ H100 GPUs, while VILA-40B is trained using ZeRO3 setting with $32$ H100 GPUs. The batch size (with accumulation) is set to $64$ and the learning rate is $1e^{-3}$. The training of Frame-Voyager is conducted over $40$ epochs requiring approximately $8$ hours for VILA-8B whereas over $20$ epochs for VILA-40B, taking around $20$ hours. All model inferences are performed on $8$ H100 GPUs.
@@ 4.2 Results and Analysis
Comparison with state-of-the-art methods. Table 1 presents the comparison of Frame-Voyager with leading Video-LLMs, organized by the size of their backbone LLMs. For models with LLMs of 8B parameters or fewer, we evaluate Video-LLaVA (Lin et al., 2024a), Qwen-VL-Chat (Bai et al., 2023), ST-LLM (Liu et al., 2024d), VideoChat2 (Li et al., 2023b), Chat-UniVi-V1.5 (Jin et al., 2024b), VideoLLaMA2 (Cheng et al., 2024), LLaVA-NeXT-QW2 (Liu et al., 2024b) and LongVILA (Xue et al., 2024). For larger models, we compare Frame-Voyager with VideoLLaMA2 (Cheng et al., 2024), VITA (Fu et al., 2024b) and LLaVA-NeXT-Video (Zhang et al., 2024c).
Among the models with LLMs of 8B parameters or fewer, Frame-Voyager achieves the best overall performance. On the Video-MME benchmark (without subtitles), it outperforms the vanilla VILA-8B by $3.0\%$, with a notable $3.6\%$ gain on long videos. Remarkably, Frame-Voyager, using only $8$ frames as input into VILA, surpasses the VILA variant LongVILA, which utilizes $128$ and $256$ frames and requires additional training and system design. LongVILA improves its performance by inputting more frames, while Frame-Voyager improves through querying more informative frame combinations, without changing the frame length limits of VILA. Frame-Voyager outperforms LongVILA even on extremely long videos (average $41$ minutes). This suggests that simply increasing the input number of frames may not always lead to better performance, since incorporating more frames might introduce noises and irrelevant information. More frames also reduce computing efficiency. Besides, among the larger Video-LLMs (with 8×\!$\times$7B and 34B LLM backbones), Frame-Voyager consistently brings notable improvements over the vanilla VILA and other state-of-the-art models. These results highlight the importance of selecting and utilizing the optimal information from video for efficient video understanding. Last but not least, VILA-8B and VILA-40B employ distinct vision encoders and LLM backbones, as outlined in Section 4.1. Thus, the consistent performance improvements indicate that Frame-Voyager functions as a plug-and-play module compatible with different Video-LLM architectures.
@@
To evaluate the effectiveness of Frame-Voyager on Video-LLMs, we conduct experiments with several baseline methods, all utilizing the same VILA-8B backbone, and the same number of frames for a fair comparison. We test rule-based shot boundary detection (SBD) methods, including Histogram (Sheena & Narayanan, 2015), Edges Change Ratio (Nasreen & Dr Shobha, 2013), and Motion (Wolf, 1996), which select frames based on significant transitions in texture, structure, and motion. Frames with the most substantial changes are chosen as extracted frames. We also include Katna [^1] https://github.com/keplerlab/katna, a cluster-based method that extracts histograms from all frames and uses K-means clustering to select most representative frames near cluster centers. In addition, two frame-text matching methods (Liang et al., 2024; Wang et al., 2024), VILA-Embedding and CLIP (Radford et al., 2021), are employed to retrieve frames by calculating cosine similarity between query inputs and individual frames, leveraging embedding layers from VILA’s LLM and CLIP. Implementation details are presented in Appendix A.
@@
[…]
@@
In Figure 3, we demonstrate the performance of Frame-Voyager on the Video-MME (without subtitles) by varying the number of chosen frames and comparing it to uniform sampling. Across different numbers of frames, Frame-Voyager consistently outperforms uniform sampling. Notably, Frame-Voyager achieves better results using only half the frames, e.g., the $8$-frame Frame-Voyager surpasses the $16$-frame uniform sampling. However, as the number of extracted frames increases, the performance gap between Frame-Voyager and uniform sampling narrows. We think this is reasonable on the current public benchmarks which usually do not require a very large number of frames to answer the query. Using more frames may introduce unnecessary information, which could even limit performance gains.
@@
In Figure 6, we present one sampled case from Video-MME (Fu et al., 2024a). The (video, query) pair, with the query “What is the small flying black dot at the start of the video”, is evaluated across different methods. The frames obtained using uniform sampling provide limited information, with irrelevant background frames included. When using CLIP for frame-query matching, the retrieved frames show a small black dot in the first frame, followed by frames related to the terms “virus” and “protein”. However, these frames do not clarify what the small black dot represents.
In contrast, our model, Frame-Voyager, effectively queries frame combinations by analyzing relationships between frames and modeling temporal information. This enables Frame-Voyager to capture the critical details at the start of the video. The selected frame combinations reveal the trajectory of the “small black dot in flight”, ultimately identifying the “dot” as a “virus”. Additional case studies are provided in Appendix 10.

### v3

@@ 1 Introduction
Recent studies (Liu et al., 2023; 2024a; Li et al., 2024c; Lin et al., 2024b) explore integrating Large Language Models (LLMs, Stiennon et al. (2020); Gao et al. (2023); OpenAI (2023); Touvron et al. (2023); Jiang et al. (2023); Yang et al. (2024)) with visual foundation models (e.g., Vision Transformer (ViT, Dosovitskiy et al. (2021)) and cross-modal projectors (Wang et al., 2018; Lin et al., 2024a; Liu et al., 2023). In this paper, we focus on Video-LLMs. Existing Video-LLMs (Zhang et al., 2023a; Cheng et al., 2024; Li et al., 2024b) usually treat the video as a sequence of image frames. The key challenge is that the entire video can not be fed into the model due to LLMs’ token length limitation (Xue et al., 2024). Meanwhile, arbitrarily increasing the token length of the model (Miao et al., 2023; Wan et al., 2024; Xiong et al., 2024; Zhang et al., 2024a) may lead to the “lost-in-the-middle” issue (Liu et al., 2024c) and introduce significant computational complexity.
To mitigate this issue, some efforts propose to select only a subset of frames as input, e.g., through uniform sampling (Wang et al., 2019; Lin et al., 2024b; Cheng et al., 2024) or text-frame matching (Liang et al., 2024; Wang et al., 2024a; Yu et al., 2024). The uniform sampling strategy evenly samples frames in videos, while text-frame matching typically retrieves a set of relevant frames by calculating semantic similarities, e.g., using CLIP (Radford et al., 2021), between the query and each frame. However, the uniform sampling fails to account for the information density variations in the videos. For instance, in the video question answering task, answering different questions may rely on distinct video segments or frames (Fu et al., 2024a). Meanwhile, text-frame matching is inadequate for complex video understanding tasks that require multi-frame or temporal reasoning, such as tracking the progression of an action or understanding cause-and-effect relationships over time. For instance, in the video summarization task, simply matching frames might overlook the subtle transitions that connect scenes, while in temporal reasoning tasks—such as answering why does the woman need to drink water at the beginning of the video?—it is crucial to concentrate on the beginning part of the video. These matching-based methods fail to account for these frame-to-frame interactions and the relative positional information essential for a comprehensive understanding of the video. To solve these problems, we introduce an innovative approach named Frame-Voyager that learns to query the subset of frames in a combinational manner, rather than retrieving individual frames separately. This capability is essential for understanding dynamic scenes and the global context of events.
@@ 2 Related Work
However, in existing Video-LLM models, a single frame is typically represented by $64$-$256$ visual tokens (Liu et al., 2023; Lin et al., 2024b; Cheng et al., 2024). Due to the input limitations of LLMs, the number of frames that can be processed by Video-LLM models is often constrained. Some studies apply techniques for handling long LLM inputs to support more frames (Miao et al., 2023; Wan et al., 2024; Xiong et al., 2024; Xue et al., 2024; Song et al., 2024), but this approach significantly increases computational complexity and can lead to issues such as the “lost-in-the-middle” effect and hallucinations (Liu et al., 2024c). Other studies, while keeping the frame input limit unchanged, use alternative sampling strategies instead of the default uniform sampling to obtain higher-quality frames as inputs. Some initial attempts focus on identifying transition frames (Lu & Grauman, 2013; Rochan et al., 2018; Rochan & Wang, 2019) or using frame clustering (Liang et al., 2024; Han et al., 2024) to find central frames, but these methods often overlook the information from the query. Subsequent research treats this problem as an individual text-frame (segment or cluster) matching task (Wang et al., 2024a; Yu et al., 2024; Wang et al., 2024b; 2025; c), attempting to select frames that are semantically closest to the query. However, this method is sub-optimal as it ignores frame-to-frame relationships, failing in complex video understanding tasks requiring multi-frame or temporal information. An alternative way is to migrate the grounded video question answering (GVQA) methods (Xiao et al., 2024; Liu et al., 2025) to general video question answering. However, they focus on identifying specific continuous temporal segments directly related to a question (Zhang et al., 2020; 2022; 2023b), which cannot meet the requirements of general video question answering.
Therefore, to the best of our knowledge, our method is the first to consider the combination of frames as a whole, aiming to find the optimal combination that can best answer the query under the constraint of frame length limitations.
@@ 3.1 Data Collection
Our pipeline is based on a simple intuition: if one frame combination is better than another, it will produce a lower language modeling loss when used as input to any trained Video-LLM. Here we adopt loss instead of correctness as a metric, since correct answers may be generated based on the Video-LLM’s inherent language prior without referring to any input content (Xiao et al., 2024). In contrast, loss reflects the model’s confidence (Guo et al., 2017; Kadavath et al., 2022) in producing the answer given the frame combination and question, where lower loss indicates higher confidence. For each (video, query) pair, we evaluate all possible combinations of $T$ frames selected from a total of $M$ video frames, resulting in $\mathcal{C}(M,T)$ combinations, where $\mathcal{C}(M,T)$ is the binomial coefficient representing the number of ways to choose $T$ items from $M$. Each combination, along with the query, is then input into a trained reference Video-LLM to calculate the combination loss, i.e., language modeling loss against the ground-truth answer. We collect the loss values for each combination and rank them from best to worst by sorting the losses in ascending order. It is worth noting that as $M$ increases, the potential number of combinations $\mathcal{C}(M,T)$ grows exponentially, making exhaustive traversal computationally infeasible. For example, when $M=64$ and $T=8$, the total number of combinations is approximately $\mathcal{C}(64,8)\approx 4\times 10^{9}$. Considering that the majority of the training data are with short videos, we use smaller combinations during training, such as $\mathcal{C}(16,2)$ or $\mathcal{C}(32,4)$. We observe from experiments that models trained with smaller combinations exhibit generalization capabilities when larger values of $M$ and $T$ are used during inference for longer video or complex reasoning. The specific choices of $M$ and $T$ employed in our experiments are detailed in Section 4.1.
@@ 3.2 Model Training and Inference
+Frame-Voyager 8B 50.5 / 53.6 60.3 / 65.0 47.3 / 50.3 43.9 / 45.3 49.8 55.7 60.8
[…]
+Frame-Voyager 7B 57.5 /    - 67.3 /    - 56.3 /    - 48.9 /    - 65.6 48.4 73.9
@@
+Frame-Voyager 34B 60.0 / 63.8 70.3 / 73.1 58.3 / 62.7 51.2 / 55.7 61.1 57.9 67.3
4 Experiments
4.1 Experiment Settings
Backbone Models. We use three variants of the state-of-the-art Video-LLM, i.e., VILA (Lin et al., 2024b): VILA-8B&40B (Wang et al., 2024b) and LLaVA-OneVision-7B (Li et al., 2024b). VILA-8B employs SigLIP (Zhai et al., 2023) as its visual encoder and Llama3-8B (Dubey et al., 2024) as its LLM backbone, while VILA-40B utilizes InternViT-6B (Chen et al., 2024) for visual encoding and Yi-34B (Young et al., 2024) as its LLM backbone.
Benchmarks. We evaluate Frame-Voyager on four widely-adopted video benchmarks: Video-MME (Fu et al., 2024a), MLVU (Zhou et al., 2024), NextQA (Xiao et al., 2021) and ActivityNet-QA (Yu et al., 2019). The former two evaluation datasets are tailored for long video assessments, while the latter two focus on short videos. We uniformly downsample the video to $128$ ($16$) candidate frames and each time select $8$ frames to compose a frame combination. The LMMs-Eval Library (Li et al., 2024a) is used for evaluation, and accuracy is reported across all benchmarks. Note that we report the accuracy scores of Video-MME under both without (no sub.) and with (sub.) subtitles settings.
@@ 4.1 Experiment Settings
Implementation Details. During the Frame-Voyager dataset construction, VILA-8B is consistently adopted as the reference Video-LLM for generating loss due to resource limitations. During training, we set $K\!=\!4$ for sampling combination ranking data. VILA-8B is trained using DeepSpeed (Aminabadi et al., 2022) ZeRO2 with $8$ H100 GPUs, while VILA-40B is trained using ZeRO3 setting with $32$ H100 GPUs. The batch size (with accumulation) is set to $64$ and the learning rate is $1e^{-3}$. The training of Frame-Voyager is conducted over $40$ epochs requiring approximately $8$ hours for VILA-8B whereas over $20$ epochs for VILA-40B, taking around $20$ hours. All model inferences are performed on $8$ H100 GPUs.
@@ 4.2 Results and Analysis
Comparison with state-of-the-art methods. Table 1 presents the comparison of Frame-Voyager with leading Video-LLMs, organized by the size of their backbone LLMs. For models with LLMs of 8B parameters or fewer, we evaluate Video-LLaVA (Lin et al., 2024a), Qwen-VL-Chat (Bai et al., 2023), ST-LLM (Liu et al., 2024d), VideoChat2 (Li et al., 2023), Chat-UniVi-V1.5 (Jin et al., 2024b), VideoLLaMA2 (Cheng et al., 2024), LLaVA-NeXT-QW2 (Liu et al., 2024b), LLaVa-OneVision (Li et al., 2024b) and LongVILA (Xue et al., 2024). For larger models, we compare Frame-Voyager with VideoLLaMA2 (Cheng et al., 2024), VITA (Fu et al., 2024b) and LLaVA-NeXT-Video (Zhang et al., 2024c).
Among the models with LLMs of 8B parameters or fewer, Frame-Voyager achieves the best overall performance. On the Video-MME benchmark (without subtitles), it outperforms the vanilla VILA-8B by $3.0\%$, with a notable $3.6\%$ gain on long videos. Remarkably, Frame-Voyager, using only $8$ frames as input into VILA, surpasses the VILA variant LongVILA, which utilizes $128$ and $256$ frames and requires additional training and system design. LongVILA improves its performance by inputting more frames, while Frame-Voyager improves through querying more informative frame combinations, without changing the frame length limits of VILA. Frame-Voyager outperforms LongVILA even on extremely long videos (average $41$ minutes). This suggests that simply increasing the input number of frames may not always lead to better performance, since incorporating more frames might introduce noises and irrelevant information. More frames also reduce computing efficiency. Besides, among the larger Video-LLMs (with 8×\!$\times$7B and 34B LLM backbones), Frame-Voyager consistently brings notable improvements over the vanilla VILA and other state-of-the-art models. These results highlight the importance of selecting and utilizing the optimal information from video for efficient video understanding. Last but not least, VILA-8B and VILA-40B employ distinct vision encoders and LLM backbones, as outlined in Section 4.1. Thus, the consistent performance improvements indicate that Frame-Voyager functions as a plug-and-play module compatible with different Video-LLM architectures. Results on additional benchmarks (MVBench (Li et al., 2024d), STAR (Wu et al., 2021), and EgoSchema (Mangalam et al., 2023)) can be found in Appendix D.
@@
To evaluate the effectiveness of Frame-Voyager on Video-LLMs, we conduct experiments with several baseline methods, all utilizing the same VILA-8B backbone and the same number of frames for a fair comparison. We test rule-based shot boundary detection (SBD) methods, including Histogram (Sheena & Narayanan, 2015), Edges Change Ratio (Nasreen & Dr Shobha, 2013), Motion (Wolf, 1996), and MDF (Han et al., 2024), which select frames based on significant transitions in texture, structure, motion and inherent similarity. Frames with the most substantial changes are chosen as extracted frames. We also include Katna [^1] https://github.com/keplerlab/katna, a cluster-based method that extracts histograms from all frames and uses K-means clustering to select most representative frames near cluster centers. In addition, six frame-text matching methods (Liang et al., 2024; Wang et al., 2024a), including VILA-Embedding, CLIP (Radford et al., 2021), SigLIP (Zhai et al., 2023), InternViT-6B (Chen et al., 2024), TempGQA (Xiao et al., 2024) and SeViLA (Yu et al., 2024), are employed to retrieve frames by calculating cosine similarity between query inputs and individual frames. Implementation details are presented in Appendix A.
@@
(a) Video-LLM Backbone: VILA-8B.
(b) Video-LLM Backbone: LLaVA-OneVision-7B.
@@
In Figure 3, we demonstrate the performance of Frame-Voyager on the Video-MME (without subtitles) by varying the number of chosen frames and comparing it to uniform sampling. Across different numbers of frames, Frame-Voyager consistently outperforms uniform sampling. Notably, Frame-Voyager achieves better results using only half the frames, e.g., the $8$-frame Frame-Voyager surpasses the $16$-frame uniform sampling. However, as the number of extracted frames increases, the performance gap between Frame-Voyager and uniform sampling narrows. There are two primary factors that influence the performance. First, public benchmarks typically do not require a large number of frames to answer queries. Incorporating more frames may introduce unnecessary information and potentially limit performance gains. Second, performance is constrained by the model’s capabilities. Given a selected frame combination, the inference model, e.g., the frozen VILA, determines the performance’s upper bound. For instance, Figure 5 demonstrates that VILA generally underperforms on OCR and counting tasks.
@@
Figure 4: RQ4. Performance of Frame-Voyager reusing different parts of VILA-8B on Video-MME (without subtitles).
Figure 5: RQ5. Performance of Frame-Voyager, CLIP, and uniform frame sampling on six question types of Video-MME.
@@
In Figure 9 of Appendix I, we present one sampled case from Video-MME (Fu et al., 2024a). The (video, query) pair, with the query “What is the small flying black dot at the start of the video”, is evaluated across different methods. The frames obtained using uniform sampling provide limited information, with irrelevant background frames included. When using CLIP for frame-query matching, the retrieved frames show a small black dot in the first frame, followed by frames related to the terms “virus” and “protein”. However, these frames do not clarify what the small black dot represents.
In contrast, our model, Frame-Voyager, effectively queries frame combinations by analyzing relationships between frames and modeling temporal information. This enables Frame-Voyager to capture the critical details at the start of the video. The selected frame combinations reveal the trajectory of the “small black dot in flight”, ultimately identifying the “dot” as a “virus”. Additional two case studies are provided in Figure 10 and Figure 11 of Appendix I.
RQ7: How does Frame-Voyager maintain efficiency and scalability when processing an increased number of selected frames?
As mentioned in Section 4.1, when training, we sample 4(2) frames from 32(16) frames, and when inference, we directly choose 8 frames from 128 frames. This demonstrates that our Frame-Voyager can be generalized from smaller numbers to larger numbers of selected frames. It is aligned with most of the existing Video-LLMs, i.e., trained primarily on short videos, yet they can also generalize to long videos during testing (Lin et al., 2024b; Li et al., 2024b). Additionally, we use two pruning methods to address efficiency concerns and reduce computational resource consumption of the data collection pipeline in Section 3.1. The results show that the data construction time can be reduced to only $4.4\%$ of the full version, while maintaining comparable performance across all benchmarks. The pruning details can be found in Appendix G.